OpenStudy (anonymous):

40 sezelle1 A population of horses has individuals that are either gray or chestnut colored. The gray color allele, G, is dominant over the chestnut color allele, g. If the frequency of the gray allele in the population is 0.8, what percent of horses are gray? A.) 64 percent B.) 96 percent C.) 32 percent D.) 80 percent

OpenStudy (anonymous):

1 = p^2 + 2pq + q^2 this is the equation you have to use

OpenStudy (anonymous):

p^2 = frequency of homozgyous dominant individuals 2pq = frequency of heterozygotes q^2 = frequency of homozgous recessive individuals.

OpenStudy (anonymous):

gray is dominant so p=0.8 and since you want to find p^2 and 2pq (because you want gray-phenotype horses) p^2 is 0.8 x 0.8 and 2pq... p+q=1 q=0.2 2pq = 2 x 0.8 x 0.2 and add your results together and you will get 0.96 so that means 96% are gray (;

OpenStudy (anonymous):

nice job thanks! follow me!!!!!!!:)